OpenStudy (anonymous):

To sirm3d #28 plz :)

OpenStudy (anonymous):

@sirm3d

OpenStudy (sirm3d):

implicit differentiation.

OpenStudy (anonymous):

for implicit, do u need to do it for both x and y

OpenStudy (sirm3d):

\[\large x^3 \cos y +e^x=1+e\]\[\large 3x^2 \cos y + x^3 (-\sin y )\frac{ dy }{ dx }+e^x =0 + 0\]

OpenStudy (sirm3d):

whenever you differentiate an expression with y, don't forget the rule to insert dy/dx

OpenStudy (anonymous):

ahh ok, let me try that out :D just need to isolate dy/dx and then sub in both x and y to solve for dy/dx

OpenStudy (sirm3d):

why don't you sub first, isolate later.

OpenStudy (anonymous):

is it easier :D ok, i will try that haha btw, did u learn statistics?

OpenStudy (anonymous):

oh yea it is easier, cuz cos (pi/2) = 0 , and ans is e :D

OpenStudy (sirm3d):

it's not my forte, but i can answer stat questions too.

OpenStudy (sirm3d):

\[\Huge \color{red} \checkmark \]

OpenStudy (anonymous):

either taking it in summer or next year :D

OpenStudy (anonymous):

I should get #29, wait LOL.. i've been doing so many trig sub..how can i not get this lol,, wait >.>

OpenStudy (sirm3d):

i'll be around. it not like the world is going to end this Dec 21, 2012. LOL

OpenStudy (anonymous):

lol i hope it ends one day earler, never need to see my final mark :D

OpenStudy (anonymous):

and u said if trig is even, u use by parts right

OpenStudy (anonymous):

odd is subs

OpenStudy (sirm3d):

that works only for sine and cosine. tangent and secant behave differently.

OpenStudy (anonymous):

how do they behave? cuz now i'm stuck with a tan ^4 (X)

OpenStudy (sirm3d):

x=3sec theta ?

OpenStudy (anonymous):

yea

OpenStudy (sirm3d):

\[\large \int\limits \frac{3 \tan \theta}{(3 \sec \theta)^3}3\sec \theta \tan \theta d \theta=(1/3)\int\limits \frac{ \tan^2 \theta }{ \sec^2 \theta } d \theta\]

OpenStudy (sirm3d):

it looks like cos^2 theta after simplification

OpenStudy (anonymous):

tan^2 theta x 1/ sec ^2 theta = tan^4 theta isnt it

OpenStudy (sirm3d):

tan = sin / cos so tan^2/sec^2 = (sin^2)/(cos^2 sec^2) = sin^2, not cos ^2

OpenStudy (sirm3d):

(cos)(sec) = 1

OpenStudy (anonymous):

\[\frac{ \sin^2 x }{ \cos^2 x } * \frac{ \sin^2 x }{ \cos^2 x }\]

OpenStudy (anonymous):

i thought it's like that..

OpenStudy (sirm3d):

1/sec = cos

OpenStudy (anonymous):

o..

OpenStudy (anonymous):

so sec = cos / sin but 1 / sec is not sin/ cos

OpenStudy (sirm3d):

sec = 1/cos, tan = sin/cos = sec/csc csc = 1/sin, cot = cos/sin = csc/sec

OpenStudy (sirm3d):

tan/sec = tan * (1 / sec) = (sin/cos)(cos)=sin

OpenStudy (anonymous):

omg. how can i get that wrong.. that trig sub confused me.. cuz i go directly from sec to tan..and now messed the easy stuff

OpenStudy (sirm3d):

sec^2 = tan^2 + 1

OpenStudy (anonymous):

my opening sentance was i must know how to do this because i have done so many trig subs my ending sentance is i have done too many trig subs, thinking tan <-> sec not tan <-> cot @_@ lol but i remember it now :D

OpenStudy (sirm3d):

the burden of remembering all the prerequisites makes calculus more difficult.

OpenStudy (anonymous):

so now since sin is even, then use sub?

OpenStudy (anonymous):

nvm..

OpenStudy (anonymous):

no derivative present..

OpenStudy (sirm3d):

even ---> half-angle odd ----> sub

OpenStudy (anonymous):

then what's by parts? even or odd?

OpenStudy (sirm3d):

you noticed no derivative. that means you're learning.

OpenStudy (sirm3d):

even ---> half-angle

OpenStudy (sirm3d):

cos^2 x = (1 + cos 2x)/2 sin^2 x = (1 - cos 2x)/2

OpenStudy (anonymous):

yea, learned alot from u :D

OpenStudy (anonymous):

\[\frac{ 1 }{ 6 }\sec ^{-1}(\frac{ x }{ 3 }) -\frac{ \sqrt{x^2-9} }{ x } \] how come i got this..

OpenStudy (sirm3d):

\[\int\limits=\frac{ \theta }{ 2 }-\frac{ \sin 2 \theta }{ 4 }+C\]

OpenStudy (sirm3d):

sin 2theta = 2 sin theta cos theta

OpenStudy (anonymous):

oh..u cant directly sub it in

OpenStudy (sirm3d):

oh, there's stilla 1/3 multiplier that i missed. just insert it later.

OpenStudy (sirm3d):

x=3sec theta, not x = 3sec 2theta.

OpenStudy (anonymous):

okay :D got the ans . i think is also save to conclude the ans is D by seeing a 1/6

OpenStudy (sirm3d):

right.

OpenStudy (anonymous):

i will try #30 first, i really think i will get it !! it's partial fraction haha

OpenStudy (sirm3d):

okay. then i'll teach you another trick in partial fractions.

OpenStudy (anonymous):

yes, plz :D

OpenStudy (anonymous):

transfer all your knowledge to me haha =P

OpenStudy (anonymous):

i got A= 55/9... do u think ths number looks a bit weird..

OpenStudy (sirm3d):

it is. its not even among the choices.

OpenStudy (anonymous):

but the A value does not necessarily reflect in the final ans isnt it?

OpenStudy (anonymous):

but yea.. A= 3

OpenStudy (anonymous):

and it does appear in the ans @@

OpenStudy (anonymous):

that means we can eliminate the option choice to D or E ?

OpenStudy (sirm3d):

now you're beating the choices.

OpenStudy (sirm3d):

C also has a 3, so you can't count it out.

OpenStudy (anonymous):

but i only have A B and C , that means i should have only 3 parts, and C and E option have 4, save to conclude it's D?

OpenStudy (sirm3d):

yes, by number of terms, it's D. BUT... the coefficient is (3/2), not 3.

OpenStudy (anonymous):

i think i just did it wrong..i got B as 0..

OpenStudy (anonymous):

oh yea, i know it cant be 0 because ther eis a (x-2)^2 < which will become x-2 after integration

OpenStudy (sirm3d):

@_@

OpenStudy (anonymous):

wait i know why..sorry lol..keep making mistakes ..

OpenStudy (anonymous):

yea ts C!

OpenStudy (anonymous):

A= 3 B=-1 C= 2

OpenStudy (anonymous):

it's 3/2 because ln2 +1 < the 2 is taken out becoming 3/2

OpenStudy (sirm3d):

Good. you spotted the mistake in choice D.

OpenStudy (anonymous):

my mistake is really stupid..i copied the question wrong lol

OpenStudy (sirm3d):

but it is still the best answer, isn't it.

OpenStudy (anonymous):

yes :D finally done my exam practice ques and practice exam, but still feel so unprepared :/

OpenStudy (sirm3d):

care to try this partial fraction problem? \[\large \int\limits \frac{7x^2-32x+31}{(x-1)(x-2)(x-3)}dx\]

OpenStudy (anonymous):

ok :D

OpenStudy (sirm3d):

or this more challenging, \[\large \int\limits \frac{2x^3+3x-1}{(x-1)^4}dx\]

OpenStudy (anonymous):

i will do all haha

OpenStudy (anonymous):

3 ln (x-1) + 5 ln (x-2) - ln (x-3)

OpenStudy (sirm3d):

solution to 1st problem is right.

OpenStudy (anonymous):

you should do (x^2 -1)^4 in the denominator lol

OpenStudy (sirm3d):

it's only (x-1)^4. hehehe.

OpenStudy (sirm3d):

let's write \[\frac{2x^3 + 3x - 1}{(x-1)^4}=\frac{ A }{ x-1 }+\frac{ B }{ (x-1)^2 }+\frac{ C }{ (x-1)^3 }+\frac{ D }{ (x-1)^4 }\]

OpenStudy (anonymous):

wait!

OpenStudy (anonymous):

I did it , wait

OpenStudy (anonymous):

i was saying u can try x^2 -1 while u are waiting for me :D

OpenStudy (anonymous):

I got 2 ln (x-1) + 3/ (x-2)^2 + 1/ (x-1)^3

OpenStudy (sirm3d):

1/(x^2 - 1) ?, there's a standard integration formula for that

OpenStudy (anonymous):

i got b as 0 again..whydo i always get B=0..

OpenStudy (anonymous):

really, cuz x^2 u need to do Ax+B thing.. which will turn into a mess

OpenStudy (sirm3d):

\[\large \int\limits \frac{ Ax+B }{ x^2-a^2 }dx=\frac{ A }{ 2 }\ln \left| x^2-a^2 \right|+\frac{ B }{ 2a }\ln \left| \frac{ x-a }{ x+a } \right|+C\]

OpenStudy (anonymous):

ohh.. i dont think i can memorize this formula for now.. but later for sure :P